# Upper Approximation Bounds for Neural Oscillators

### v1

@@ A.1 Proof of Lemma 1
$\begin{array}[]{l}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}(\omega)=\int_{-\infty}^{+\infty}{{e^{-{\rm{i}}\omega\tau}}{{{\tilde{\mathbf{f}}}}_{t}}(\tau){\rm{d}}\tau}=-{\rm{i}}\int_{-t}^{t}{\sin(\omega\tau){{{\tilde{\mathbf{f}}}}_{t}}(\tau){\rm{d}}\tau}\\ \;\;\;\;\;\;\;\;\;\;\;\;=-2{\rm{i}}\int_{0}^{t}{\sin(\omega\tau){{\bf{f}}_{t}}(\tau){\rm{d}}\tau}=-2{\rm{i}}\int_{0}^{t}{\sin(\omega\tau)\left[{{\bf{u}}(t-\tau)-{\bf{u}}(0)}\right]{\rm{d}}\tau}\\ \;\;\;\;\;\;\;\;\;\;\;\;=-2{\rm{i}}\left\{{{\mathcal{L}_{t}}{\bf{u}}(\omega)+\frac{{{\bf{u}}(0)}}{\omega}\left[{\cos(\omega t)-1}\right]}\right\}.\end{array}$ (40)
@@
$\begin{array}[]{l}\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-\left[{{\bf{u}}(t-\tau)-{\bf{u}}(0)}\right]}\right|\\ \;\;\;\;\;=\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right|=\left|{\int_{-v}^{0}{\left[{{{{\tilde{\mathbf{f}}}}_{t}}(\tau-\lambda)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right]{\rho_{v}}(\lambda){\rm{d}}\lambda}}\right|\\ \;\;\;\;\;\leq\int_{-v}^{0}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\tau-\lambda)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right|{\rho_{v}}(\lambda){\rm{d}}\lambda}\mathop{\leq}\limits^{(a)}\int_{-v}^{0}{\left|{\mathbf{\upphi}_{K}(v)}\right|{\rho_{v}}(\lambda){\rm{d}}\lambda}=\left|{\mathbf{\upphi}_{K}(v)}\right|,\end{array}$ (43)
@@
$\tilde{\mathbf{f}}_{v}(\tau)$, $\tilde{\mathbf{f}}_{t}(\tau)$, and $\rho_{v}(\tau)$ are compactly supported over $\tau\in[-2T,2T]$ for all $t\in[0,T]$ and all $v\in\left(0,T\right)$. $\tilde{\mathbf{f}}_{v}(\tau)$ and $\rho_{v}(\tau)$ are continuous over $\tau\in\mathbb{R}$. $\tilde{\mathbf{f}}_{v}(\tau)$ is odd with respect to $\tau=-0.5v$ because $\tilde{\mathbf{f}}_{t}(\tau)$ is an odd function and $\rho_{v}(\tau)$ is symmetric with respect to $\tau=-0.5v$, thereby $\int_{-\infty}^{+\infty}\tilde{\mathbf{f}}_{v}(\tau)\mathrm{d}\tau=\int_{-2T}^{2T}\tilde{\mathbf{f}}_{v}(\tau)\mathrm{d}\tau=\mathbf{0}$. Then, $\tilde{\mathbf{f}}_{v}(\tau)$ over $\tau\in[-2T,2T]$ can be expressed using the Fourier series
${{\tilde{\mathbf{f}}}_{v}}(\tau)=2\sum\limits_{n=1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{S}{{{\tilde{\mathbf{f}}}}_{v}}({\omega_{n}})}\right]}\mathop{=}\limits^{(a)}\frac{1}{{2T}}\sum\limits_{n=1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]},$ (44)
where $\omega_{n}=\pi n/{(2T)}$, $\mathcal{S}{{\tilde{\mathbf{f}}}}_{v}(\omega_{n})=1/(4T)\int_{-2T}^{2T}e^{-\mathrm{i}\omega_{n}}\tilde{\mathbf{f}}_{v}(\tau){\mathrm{d}\tau}=\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})/(4T)$ is the Fourier coefficient of $\tilde{\mathbf{f}}_{v}(\tau)$, Step (a) is valid because $\tilde{\mathbf{f}}_{v}(\tau)$ is the convolution between $\tilde{\mathbf{f}}_{t}(\tau)$ and $\rho_{v}(\tau)$, and $\mathcal{H}\rho_{v}(\omega)$ is the Fourier transform of $\rho_{v}(\tau)$. $\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})/(2T)$ in Eq. (44) can be bounded by
$\begin{array}[]{l}\frac{1}{{{2T}}}\left|{\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right|\\ \;\;\;\;\;=\frac{1}{{{2T}}}\left|{\int_{-{2T}}^{{2T}}{{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}}}_{v}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;\;\;=\frac{1}{{{2T}}}\left|{\left.{\frac{1}{{-{\rm{i}}{\omega_{n}}}}{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}}}_{v}}(\tau)}\right|_{-{2T}}^{{2T}}-\frac{1}{{-{\rm{i}}{\omega_{n}}}}\int_{-{2T}}^{{2T}}{{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}^{\prime}}}_{v}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;\;\;=\frac{1}{{{2T}}}\left|{\left.{-\frac{1}{{-\omega_{n}^{2}}}{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}^{\prime}}}_{v}}(\tau)}\right|_{-{2T}}^{{2T}}-\frac{1}{{\omega_{n}^{2}}}\int_{-{2T}}^{{2T}}{{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}^{\prime\prime}}}_{v}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;\;\;\leq\frac{{{2T}}}{{{\pi^{2}}{n^{2}}}}\int_{-{2T}}^{{2T}}{\left|{{{{\tilde{\mathbf{f}}^{\prime\prime}}}_{v}}(\tau)}\right|{\rm{d}}\tau}\leq\frac{{8T^{2}}}{{{\pi^{2}}{n^{2}}}}\mathop{\max}\limits_{\tau\in[-{2T},{2T}]}\left|{{{{\tilde{\mathbf{f}}^{\prime\prime}}}_{v}}(\tau)}\right|.\end{array}$ (45)
@@
In Eq. (45), since $\rho_{v}(\tau)$ has arbitrary finite-order continuous derivatives and $\tilde{\mathbf{f}}_{v}(\tau)$, $\tilde{\mathbf{f}}_{t}(\tau)$, and $\rho_{v}(\tau)$ are compactly supported over $\tau\in[-2T,2T]$ for all $t\in[0,T]$, the boundary conditions ${\tilde{\mathbf{f}}}_{v}(-2T)={\tilde{\mathbf{f}}}_{v}(2T)={\tilde{\mathbf{f}}}_{v}^{\prime}(-2T)={\tilde{\mathbf{f}}}_{v}^{\prime}(2T)=\mathbf{0}$ are employed. ${\tilde{\mathbf{f}}}_{v}^{\prime\prime}(\tau)$ in Eq. (45) over $\tau\in[-2T,2T]$ for all $t\in[0,T]$ is bounded by
@@
$\begin{array}[]{l}\left|{{{\tilde{\mathbf{f}}^{\prime\prime}}_{v}}(\tau)}\right|=\left|{\frac{{{{\rm{d}}^{2}}}}{{{\rm{d}}{\tau^{2}}}}\int_{-\infty}^{+\infty}{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda){\rho_{v}}(\tau-\lambda){\rm{d}}\lambda}}\right|\leq\int_{-\infty}^{+\infty}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda)}\right|\left|{\frac{{{{\rm{d}}^{2}}}}{{{\rm{d}}{\tau^{2}}}}{\rho_{v}}(\tau-\lambda)}\right|{\rm{d}}\lambda}\\ \;\;\;\;\;\;\;\;\;\;\leq\int_{\tau}^{\tau+v}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}_{v}^{\prime\prime}}(t)}\right|{\rm{d}}\lambda}\leq\int_{0}^{v}{\mathop{\max}\limits_{t\in[0,T]}\left|{{\bf{u}}(t)-{\bf{u}}(0)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}_{v}^{\prime\prime}}(t)}\right|{\rm{d}}\lambda}\\ \;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}v\left|{\mathbf{\upphi}_{K}(T)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}_{v}^{\prime\prime}}(t)}\right|\mathop{\leq}\limits^{(b)}\left|{\mathbf{\upphi}_{K}(T)}\right|\frac{{{c_{\rho,2}}}}{{{v^{2}}}},\end{array}$ (46)
@@
where Step (a) employs $|\mathbf{u}(t)-\mathbf{u}(0)|\leq\left|{\mathbf{\upphi}}_{K}(t)\right|\leq\left|{\mathbf{\upphi}}_{K}(T)\right|$, Step (b) holds from $\max_{\tau\in[-v,0]}\left|{\rho}_{v}^{\prime\prime}(\tau)\right|$ $\leq c_{\rho,2}/v^{3}$, and $c_{\rho,2}=\max_{\tau\in[-1,0]}\left|{\rho}_{1}^{\prime\prime}(\tau)\right|$ is an independent constant. The bound of $|{\tilde{\mathbf{f}}}_{v}^{\prime\prime}(\tau)|$ in Eq. (46) is independent of t and $\tau$. Then, based on Eqs. (44)-(46), for every positive integer M larger than one, it can be obtained
@@
$\begin{array}[]{l}\left|{{{{\tilde{\bf f}}}_{v}}(\tau)-\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\bf f}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{\nu}}({\omega_{n}})}\right]}}\right|\\ \;\;\;=\left|{\frac{1}{{2T}}\sum\limits_{n=M+1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\bf f}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{\nu}}({\omega_{n}})}\right]}}\right|\\ \;\;\;\leq\frac{1}{{2T}}\sum\limits_{n=M+1}^{+\infty}{\left|{\mathcal{H}{{{\tilde{\bf f}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{\nu}}({\omega_{n}})}\right|}\mathop{\leq}\limits^{(a)}\sum\limits_{n=M+1}^{+\infty}{\frac{{8{T^{2}}}}{{{\pi^{2}}{n^{2}}}}\mathop{\max}\limits_{\tau\in[-{2T},{2T}]}\left|{{{{\tilde{\mathbf{f}}}^{\prime\prime}}_{v}}(\tau)}\right|}\\ \;\;\;\leq\frac{8T^{2}}{{{\pi^{2}}}}{\mathop{\max}\limits_{\tau\in[-2T,2T]}\left|{{{{\tilde{\mathbf{f}}}^{\prime\prime}}_{v}}(\tau)}\right|}\sum\limits_{n=M+1}^{+\infty}{\frac{1}{{{n^{2}}}}}\mathop{\leq}\limits^{(b)}\frac{{8T^{2}\left|{\mathbf{\upphi}}_{K}(T)\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}}}\sum\limits_{n=M+1}^{+\infty}{\frac{1}{{{n^{2}}}}}\mathop{\leq}\limits^{(c)}\frac{{8T^{2}\left|{\mathbf{\upphi}}_{K}(T)\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}},\end{array}$ (47)
where Steps (a) and (b) are from Eqs. (45) and (46), respectively, and Step (c) is derived from the integral test for convergence, see Section 3.3 in (Knopp, 1956),
$\sum\limits_{n=M+1}^{+\infty}{\frac{1}{{{n^{2}}}}}\leq\sum\limits_{m=0}^{+\infty}{\int_{M+m}^{M+m+1}{\frac{1}{{{x^{2}}}}{\rm{d}}x}}=\int_{M}^{+\infty}{\frac{1}{{{x^{2}}}}{\rm{d}}x}=\frac{1}{M},\;\;m,n\in\mathbb{Z}.$ (48)
Finally, based on Eqs. (43) and (47), for every positive integer M larger than one, $\mathbf{u}(t-\tau)$ over $\tau\in[0,t]$ satisfies
@@
$\begin{array}[]{l}\left|{{\bf{u}}(t-\tau)-\left\{{\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}+{\bf{u}}(0)}\right\}}\right|\\ \;\;\;\;=\left|{{\bf{u}}(t-\tau)-{\bf{u}}(0)-\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}}\right|\\ \;\;\;\;\leq\left|{{\bf{u}}(t-\tau)-{\bf{u}}(0)-{{{\tilde{\mathbf{f}}}}_{v}}(\tau)}\right|+\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}}\right|\\ \;\;\;\;\leq\left|{\mathbf{\upphi}_{K}(v)}\right|+\frac{{8T^{2}\left|{\mathbf{\upphi}}_{K}(T)\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}}\end{array}$ (49)
for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$. Based on Eq. (40), $1/(2T){\mathrm{Re}}[e^{\mathrm{i}\omega_{n}\tau}\mathcal{H}\tilde{\mathbf{f}}_{t}(\omega_{n})\mathcal{H}\rho_{v}(\omega_{n})]$ in Eq. (49) can be expressed as
@@
$\begin{array}[]{l}\frac{1}{{2T}}{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\\ \;\;\;\;\;=\frac{1}{T}\left\{{{\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\left\{{{\mathop{\rm Re}\nolimits}\left[{\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\sin({\omega_{n}}\tau)+{\mathop{\rm Im}\nolimits}\left[{\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\cos({\omega_{n}}\tau)}\right\}\\ \;\;\;\;\;={\alpha_{n}}\left\{{{\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin({\omega_{n}}\tau-{\theta_{n}}),\end{array}$ (50)

### v2

@@ A.1 Proof of Lemma 1
$\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}(\omega)=\int_{-\infty}^{+\infty}{{e^{-{\rm{i}}\omega\tau}}{{{\tilde{\mathbf{f}}}}_{t}}(\tau){\rm{d}}\tau}=-{\rm{i}}\int_{-t}^{t}{\sin(\omega\tau){{{\tilde{\mathbf{f}}}}_{t}}(\tau){\rm{d}}\tau}$ (45)
$\;\;\;\;\;\;\;\;\;\;\;=-2{\rm{i}}\int_{0}^{t}{\sin(\omega\tau){{\bf{f}}_{t}}(\tau){\rm{d}}\tau}=-2{\rm{i}}\int_{0}^{t}{\sin(\omega\tau)\left[{{\bf{u}}(t-\tau)-{\bf{u}}(0)}\right]{\rm{d}}\tau}$
$\;\;\;\;\;\;\;\;\;\;\;=-2{\rm{i}}\left\{{{\mathcal{L}_{t}}{\bf{u}}(\omega)+\frac{{{\bf{u}}(0)}}{\omega}\left[{\cos(\omega t)-1}\right]}\right\}.$
@@
$\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-\left[{{\bf{u}}(t-\tau)-{\bf{u}}(0)}\right]}\right|=\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right|=\left|{\int_{-v}^{0}{\left[{{{{\tilde{\mathbf{f}}}}_{t}}(\tau-\lambda)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right]{\rho_{v}}(\lambda){\rm{d}}\lambda}}\right|$ (48)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq\int_{-v}^{0}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\tau-\lambda)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right|{\rho_{v}}(\lambda){\rm{d}}\lambda}\mathop{\leq}\limits^{(a)}\int_{-v}^{0}{\left|{\mathbf{\upphi}_{K}(v)}\right|{\rho_{v}}(\lambda){\rm{d}}\lambda}=\left|{\mathbf{\upphi}_{K}(v)}\right|,$
@@
Since the compactly supported $\rho_{v}(\tau)$ belongs to a Gevrey class of order 2 (Rodino, 1993) and it has arbitrary finite-order continuous derivatives, the $r^{\text{th}}$-order derivative ${\rm{d}}^{r}{\tilde{\mathbf{f}}}_{v}(\tau)/{\rm{d}}\tau^{r}$ of ${\tilde{\mathbf{f}}}_{v}(\tau)$ over $\tau\in[-2T,2T]$ for all $t\in[0,T]$ and all positive $r\in\mathbb{Z}$ is bounded by
$\left|\frac{{\rm{d}}^{r}{\tilde{\mathbf{f}}}_{v}(\tau)}{{\rm{d}}\tau^{r}}\right|=\left|{\frac{{\rm{d}}^{r}}{{\rm{d}}\tau^{r}}\int_{-\infty}^{+\infty}{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda){\rho_{v}}(\tau-\lambda){\rm{d}}\lambda}}\right|\leq\int_{-\infty}^{+\infty}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda)}\right|\left|{\frac{{\rm{d}}^{r}}{{\rm{d}}\tau^{r}}{\rho_{v}}(\tau-\lambda)}\right|{\rm{d}}\lambda}$ (49)
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\leq\int_{\tau}^{\tau+v}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{\frac{{\rm{d}}^{r}}{{\rm{d}}\tau^{r}}{{\rho}_{v}}(t)}\right|{\rm{d}}\lambda}\leq\int_{0}^{v}{\mathop{\max}\limits_{t\in[0,T]}\left|{{\bf{u}}(t)-{\bf{u}}(0)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{\frac{{\rm{d}}^{r}}{{\rm{d}}\tau^{r}}{{\rho}_{v}}(t)}\right|{\rm{d}}\lambda}$
$\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}v\left|{\mathbf{\upphi}_{K}(T)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{\frac{{\rm{d}}^{r}}{{\rm{d}}\tau^{r}}{{\rho}_{v}}(t)}\right|\mathop{\leq}\limits^{(b)}\left|{\mathbf{\upphi}_{K}(T)}\right|\frac{c_{\rho}^{r+1}\left(r!\right)^{2}}{v^{r}},$
@@
where Step (a) employs $|\mathbf{u}(t)-\mathbf{u}(0)|\leq\left|{\mathbf{\upphi}}_{K}(t)\right|\leq\left|{\mathbf{\upphi}}_{K}(T)\right|$, Step (b) holds becasue $\rho_{v}(\tau)$ belongs to a Gevrey class of order 2 (Rodino, 1993), and $c_{\rho}>1$ is an independent constant. The bound of ${\rm{d}}^{r}{\tilde{\mathbf{f}}}_{v}(\tau)/{\rm{d}}\tau^{r}$ in Eq. (49) is independent of t and $\tau$.
$\tilde{\mathbf{f}}_{v}(\tau)$, $\tilde{\mathbf{f}}_{t}(\tau)$, and $\rho_{v}(\tau)$ are compactly supported over $\tau\in[-2T,2T]$ for all $t\in[0,T]$ and all $v\in\left(0,T\right)$. $\tilde{\mathbf{f}}_{v}(\tau)$ is odd with respect to $\tau=-0.5v$ because $\tilde{\mathbf{f}}_{t}(\tau)$ is an odd function and $\rho_{v}(\tau)$ is symmetric with respect to $\tau=-0.5v$, thereby $\int_{-\infty}^{+\infty}\tilde{\mathbf{f}}_{v}(\tau)\mathrm{d}\tau=\int_{-2T}^{2T}\tilde{\mathbf{f}}_{v}(\tau)\mathrm{d}\tau=\mathbf{0}$. Then, $\tilde{\mathbf{f}}_{v}(\tau)$ over $\tau\in[-2T,2T]$ can be expressed as
@@
${{\tilde{\mathbf{f}}}_{v}}(\tau)=\frac{1}{2T}\sum\limits_{n=1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{v}}({\omega_{n}})}\right]}\mathop{=}\limits^{(a)}\frac{1}{2T}\sum\limits_{n=1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]},$ (50)
where $\omega_{n}=\pi n/{(2T)}$, $\mathcal{H}{{\tilde{\mathbf{f}}}}_{v}(\omega_{n})$ is the Fourier transform of $\tilde{\mathbf{f}}_{v}(\tau)$
$\mathcal{H}{{\tilde{\mathbf{f}}}}_{v}(\omega_{n})=\int_{-2T}^{2T}e^{-\mathrm{i}\omega_{n}\tau}\tilde{\mathbf{f}}_{v}(\tau){\mathrm{d}\tau},$ (51)
$\mathcal{H}\rho_{v}(\omega_{n})$ is the Fourier transform of $\rho_{v}(\tau)$
$\mathcal{H}\rho_{v}(\omega_{n})=\int_{-2T}^{2T}e^{-\mathrm{i}\omega_{n}\tau}\rho_{v}(\tau){\mathrm{d}\tau},$ (52)
and Step (a) is valid because $\tilde{\mathbf{f}}_{v}(\tau)$ is the convolution between $\tilde{\mathbf{f}}_{t}(\tau)$ and $\rho_{v}(\tau)$.
Eq. (49) indicates that $\left|{\mathbf{\upphi}_{K}(T)}\right|c_{\rho}^{r+1}\left(r!\right)^{2}v^{-r}$ is a Lipschitz constant of ${\rm{d}}^{r-1}{\tilde{\mathbf{f}}}_{v}(\tau)/{\rm{d}}\tau^{r-1}$. Then, according to Corollary III and Theorem X of Jackson (1930), for every positive integer M larger than two and every positive integer $r$, it can be obtained
@@
$\left|{{{\tilde{\bf f}}}_{v}}(\tau)-\frac{1}{2T}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}\right|\leq\frac{2^{r}c_{F}^{r}T^{r}\left|{\mathbf{\upphi}_{K}(T)}\right|c_{\rho}^{r+1}\left(r!\right)^{2}\ln{M}}{\pi^{r}v^{r}M^{r}},$ (53)
where $c_{F}>1$ is an independent constant.
Finally, based on Eqs. (48) and (53), for every positive integer M larger than two and every positive integer $r$, $\mathbf{u}(t-\tau)$ over $\tau\in[0,t]$ satisfies
@@
$\left|{{\bf{u}}(t-\tau)-\left\{{\frac{1}{2T}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}+{\bf{u}}(0)}\right\}}\right|$ (54)
$\;\;\;\;=\left|{{\bf{u}}(t-\tau)-{\bf{u}}(0)-\frac{1}{2T}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}}\right|$
$\;\;\;\;\leq\left|{{\bf{u}}(t-\tau)-{\bf{u}}(0)-{{{\tilde{\mathbf{f}}}}_{v}}(\tau)}\right|+\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-\frac{1}{2T}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}}\right|$
$\;\;\;\;\leq\left|{\mathbf{\upphi}_{K}(v)}\right|+\frac{c_{K}^{r+1}T^{r}\left|{\mathbf{\upphi}_{K}(T)}\right|\left(r!\right)^{2}\ln{M}}{v^{r}M^{r}}$
for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$, where $c_{K}=c_{F}c_{\rho}>1$ is an independent constant. Based on Eq. (45), $\left(2T\right)^{-1}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}$ in Eq. (54) can be expressed as
$\frac{1}{2T}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}$ (55)
$\;\;\;\;\;=\frac{1}{T}\left\{{{\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\left\{{{\mathop{\rm Re}\nolimits}\left[{\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\sin({\omega_{n}}\tau)+{\mathop{\rm Im}\nolimits}\left[{\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\cos({\omega_{n}}\tau)}\right\}$
$\;\;\;\;\;={\alpha_{n}}\left\{{{\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin({\omega_{n}}\tau-{\theta_{n}}),$